OpenStudy (anonymous):

Find the perimeter of the square. A. 64 ft B. 32 ft C. 16 ft D. 8 ft

OpenStudy (anonymous):

where's the square?

OpenStudy (jordan123321):

what do you think

OpenStudy (anonymous):

here it is

OpenStudy (welshfella):

4 * 8

OpenStudy (jordan123321):

well p = 4 a

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

8 x 4 that's all you have to do

OpenStudy (jordan123321):

yes

OpenStudy (welshfella):

- as there are 4 sides all 8 ft long

OpenStudy (anonymous):

to get the perimeter

OpenStudy (jordan123321):

p = 32

OpenStudy (jordan123321):

b would b your answer

OpenStudy (ahsome):

Perimeter of square =\(4s\), where S is the side So, since the length of the side is \(8\), we can solve for the perimeter \[\Huge{Perimeter=4s}\]\[\Huge{Perimeter=4*8}\]\[\Huge{Perimeter=32m}\] Therefore, the perimeter is 32m, so B

OpenStudy (anonymous):

So b is your answer

OpenStudy (jordan123321):

yes

OpenStudy (anonymous):

formula to find perimeter of a square= 4*s (s means sides) so it's 4*8=p (p means perimeter)